# Combinatorial Selection with Costly Information

### v1

@@ 3.3 Water Filling and Surrogate Costs for General MDPs
We note that the lemma is standard (see for example (Strassen, 1965; Föllmer and Schied, 2016)) but we provide a constructive proof in Appendix A for the sake of intuition and completeness. We apply Lemma 3.5 to all tuples $(W^{*}_{\mathcal{M}},\hat{Z}_{j})$ to obtain mappings $m_{j}(\cdot)$.
We are finally ready to define the amortized cost function $\rho$ and the cost sharing vector $b$. Fix some $t\in T$, and let $j\in[k]$ and $s\in S$ be the indices of the unique action $a_{j}$ and state $s\in R_{j}$ such that $t\in T(s)$. We define
$\rho(t):=m_{j}(\max\{g_{j},\rho_{s}(t)\})$
@@
$b_{\sigma t}:=\operatorname{E}\left[\rho(t)\right]-\operatorname{E}\left[\rho_{s}(t)\right]$
and for all other $s^{\prime}\in S\setminus\{\sigma\}$ with $t\in T(s^{\prime})$, we use the same cost share $b_{s^{\prime}t}=b^{s}_{s^{\prime}t}$ that was used in $\mathcal{M}_{s}$.
We will now show that these costs satisfy the properties claimed in the lemma. Fix any committing policy $\pi\in\mathcal{C}\left(\mathcal{M}\right)$. Without loss, we will assume that $\pi$ deterministically commits to a fixed action $a_{j}$ at $\sigma$; once our results are established for these committing policies, they immediately extend to arbitrary committing policies by linearity of expectation.
  • Action independence: Drawing a terminal node $t\sim p_{\pi}$ is equivalent to first drawing a state $s\in R_{j}$ from the distribution $\mathcal{D}(\sigma,a_{j},\cdot)$ and then drawing a terminal node from the restriction of $\pi$ to $\mathcal{M}_{s}$. Consider drawing $t$ in this manner and then sampling from the distribution $\rho_{s}(t)$. By the induction hypothesis and the definition of $Z_{j}$, this provides us with a sample drawn from $Z_{j}$. Then, $\max\{g_{j},\rho_{s}(t)\}$ with $t$ drawn in this manner corresponds to an instantiation of $\hat{Z}_{j}$, and $m_{j}$ applied to that instantiation results in an instantiation of $W^{*}_{\mathcal{M}}$ by the definition of $m_{j}$ and property (1) in Lemma 3.5.
  • Cost sharing: This holds trivially by the definition of the cost shares in the starting state $\sigma$, the fact that we don’t change the cost shares in any other state $s\neq\sigma$ and the induction hypothesis.
  • Cost dominance: By the induction hypothesis, the inequality holds for all nodes other than $\sigma$. Note that $p_{\pi}$ only places non-zero probability mass on states downstream from action $a_{j}$. Denote by $\pi_{|s}$ the policy $\pi$ confined to the subprocess $\mu_{s}$ for some $s\in R_{j}$. Then we can write the total cost share paid to $\sigma$ under some policy $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ as:
    $\sum_{t\in T_{\pi}}p_{\pi}(t)b_{\sigma t}$ $=\sum_{s\in R_{j},t\in T_{s}}p_{\pi}(t)(\operatorname{E}\left[\rho(t)\right]-\operatorname{E}\left[\rho_{s}(t)\right])$
    $=\sum_{t\in T_{\pi}}p_{\pi}(t)\operatorname{E}\left[\rho(t)\right]-\sum_{s\in R_{j}}\mathcal{D}(\sigma,a_{j},s)\sum_{t\in T_{\pi_{|s}}}p_{\pi_{|s}}(t)\operatorname{E}\left[\rho_{s}(t)\right]$
@@
    $\leq\operatorname{E}\left[\hat{Z}_{j}\right]-\operatorname{E}\left[Z_{j}\right]=c(a_{j})$
    which proves the claim by noting that $p_{\pi}(\sigma)=1$. Here the first equation follows from the definition of $b_{st}$; the second just rewrites the terms separately; the third is by the action independence of $\rho$ as proved above, and by the induction hypothesis applied to $\rho_{s}$ similarly; the fourth uses property 2. in Lemma 3.5 for the first term and the definition of $Z_{j}$ for the second term; and the last equality follows from the definition of $\hat{Z}_{j}$.
@@ 4 Local Approximation and Composition Theorems
In Section 3 we showed that water filling surrogate costs provide a lower bound on the optimal cost for any instance $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$ of the matroid-min-CICS. Furthermore, when all of the MDPs $\mathcal{M}_{i}$ in the instance are Markov chains, not only is this lower bound exact but it also suggests a simple optimal policy. We will now put these concepts together to enable approximations for general instances. Following previous work on Pandora’s Box, our approach finds a committing policy $\pi_{i}\in\mathcal{C}\left(\mathcal{M}_{i}\right)$ for each MDP $\mathcal{M}_{i}$, that exhibits some notion of approximation in the local game for the MDP. Our goal is to then combine these local guarantees into a global approximation guarantee.
A natural first attempt is to find a committing policy $\pi$ for $\mathcal{M}$ that achieves approximation through optimality curves: $f_{\mathcal{M}^{\pi}}(y)\leq\alpha\cdot f_{\mathcal{M}}(y)$ for all $y$ and some fixed factor $\alpha\geq 1$. As we showed in the introduction, however, this condition is not sufficient for a global guarantee. We adopt the following strengthening from Scully and Doval (2024).
@@
Let $\mathcal{M}$ be any MDP. We say that a committing policy $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ $\alpha$-locally approximates $\mathcal{M}$ for some $\alpha\geq 1$ if
$f_{\mathcal{M}^{\pi}}(\alpha y)\leq\alpha\cdot f_{\mathcal{M}}(y)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall y\in\real.$
We say that a decision process $\mathcal{M}$ admits an $\alpha$-local approximation if there exists a committing policy $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ that $\alpha$-locally approximates it.
Local approximation has an intuitive interpretation. From the inherent connection between optimality curves and water filling surrogate costs of Definition 6, we can equivalently state the local approximation condition as
$\operatorname{E}\left[\min\{y,W^{*}_{\mathcal{M}^{\pi}}\}\right]\leq\operatorname{E}\left[\min\{y,\alpha\cdot W^{*}_{\mathcal{M}}\}\right]\;\;\forall y\in\real.$
By definition, water filling surrogate costs scale with the parameters of the underlying MDP; in other words, the surrogate cost of a scaled-up version of $\mathcal{M}$, where all action costs and values are multiplied by $\alpha$, will be precisely $\alpha\cdot W^{*}_{\mathcal{M}}$; we use $\alpha\mathcal{M}$ to denote this scaled-up version of $\mathcal{M}$. With this in mind, the local approximation condition states that an optimal policy for the local game $(\mathcal{M},y)$ under commitment $\pi$ performs better than the optimal adaptive policy in the local game $(\alpha\mathcal{M},y)$. Equivalently, the surrogate costs $\alpha W^{*}_{\mathcal{M}}$ second-order stochastically dominate the surrogate costs $W^{*}_{\mathcal{M}_{\pi}}$.
We can also establish local approximation through a stronger but conceptually easier relationship between the surrogate costs: namely, that $\alpha W^{*}_{\mathcal{M}}$ first-order stochastically dominates $W^{*}_{\mathcal{M}_{\pi}}$. Formally, we define the following alternate notion of approximation, where for a quantile $q\in[0,1]$ and a random variable $X$, we let $X(q)$ denote the $q$th quantile value of the random variable: $X(q)=\inf\{x:Pr[X\leq x]\geq q\}$.
Definition 8 (Pointwise Approximation of Surrogate Costs).
Let $\mathcal{M}$ be any MDP. We say that a committing policy $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ $\alpha$-pointwise approximates $\mathcal{M}$ for some $\alpha\geq 1$ if we have $W^{*}_{\mathcal{M}^{\pi}}(q)\leq\alpha\cdot W^{*}_{\mathcal{M}}(q)$ for all $q\in[0,1]$.
The following is immediate, but the converse is not always true.
Fact 2.
For any MDP $\mathcal{M}$, any policy that $\alpha$-pointwise approximates $\mathcal{M}$ will also $\alpha$-locally approximate $\mathcal{M}$.
Composition of Local Approximation
It is worth noting that in the special case of $\alpha=1$, local approximation coincides with Whittle’s condition (Whittle, 1980; Glazebrook, 1982), according to which an MDP $\mathcal{M}$ admits a fixed sequence of actions that is always optimal for the local game $(\mathcal{M},y)$. In that sense, local approximation can also be viewed as a relaxation of Whittle’s condition. It is well-known that if all the MDPs satisfy Whittle’s condition, then we can efficiently find an optimal solution to matroid-min-CICS. In other words, if each $\mathcal{M}_{i}$ admits a $1$-local approximation under some local commitment $\pi_{i}\in\mathcal{C}\left(\mathcal{M}_{i}\right)$, then we can deduce that the global commitment $\Pi=(\pi_{1},\cdots,\pi_{n})$ allows for a $1$-approximation with respect to the optimal adaptive policy. We refer to such statements as composition results, since they allow us to combine local guarantees on the underlying MDPs into global approximation guarantees for matroid-min-CICS.
The main strength of local approximation lies in the fact that much like Whittle’s condition, it satisfies the following composition result. The proof is straightforward from our discussion above; see Appendix B.
@@
Let $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$ be any instance of matroid-min-CICS, where each MDP $\mathcal{M}_{i}$ admits an $\alpha$-local approximation under some committing policy $\pi_{i}\in\mathcal{C}\left(\mathcal{M}_{i}\right)$. Then,
@@
$\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}^{\pi_{i}}}\right]\leq\alpha\cdot\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}\right].$
Observe that Theorem F.2 directly implies a way to efficiently approximate the optimal adaptive policy for any instance of matroid-min-CICS, assuming that the local approximation guarantees are met. We state this approach in Algorithm 1.
1
Input : An instance $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$ of matroid-min-CICS.
A tuple of committing policies $\Pi=(\pi_{1},\cdots,\pi_{n})$ with $\pi_{i}\in\mathcal{C}\left(\mathcal{M}_{i}\right)$.
2 For each MDP $\mathcal{M}_{i}$ and state $s_{i}\in S_{i}$, sample a unique action $\hat{a}_{i}(s_{i})\sim\pi_{i}(s_{i})$.
3 For all $i\in[n]$, construct the Markov chain $\mathcal{M}^{\prime}_{i}$ from $\mathcal{M}_{i}$ with only actions $\hat{a}_{i}(s_{i})$ available.
4 For all $i\in[n]$, compute the water filling indices of all the states in $\mathcal{M}^{\prime}_{i}$.
Run the water filling index policy on instance $(\mathcal{M}^{\prime}_{1},\cdots,\mathcal{M}^{\prime}_{n},\mathcal{F})$.
Algorithm 1 Local Approximation Composition Algorithm
Clearly, all the steps of the algorithm run in polynomial time to the size of the MDPs. By combining the optimality of the water filling index policy on Markov chains (Theorem 3.2) with our lower bound on the optimal adaptive policy (Theorem 3.4) and our composition theorem (Theorem 4.1), we immediately obtain the following:
Corollary 4.2.
Let $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$ be any instance of matroid-min-CICS. Instantiating Algorithm 1 with a set of committing policies $\pi_{i}$ that $\alpha$-locally approximate the MDPs $\mathcal{M}_{i}$ results to an $\alpha$-approximation to the optimal adaptive policy for $\mathbb{I}$.

### v2

@@ 3.3 Water Filling and Surrogate Costs for General MDPs
We note that the lemma is standard (see for example (Strassen, 1965; Föllmer and Schied, 2016)) but we provide a constructive proof in Appendix D for the sake of intuition and completeness. We apply Lemma 3.6 to all tuples $(W^{*}_{\mathcal{M}},\hat{Z}_{j})$ to obtain mappings $m_{j}(\cdot)$.
We are finally ready to define the amortized cost functions $\rho^{\pi}$ and the cost sharing vectors $b^{\pi}$. Fix any deterministic commitment $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ and let $j=\pi(\sigma)\in[k]$ be the fixed action that $\pi$ takes at state $\sigma$. For each state $s\in R_{j}$, we use $\pi|s\in\mathcal{C}\left(\mathcal{M}_{s}\right)$ to denote the commitment $\pi$ after we transition to state $s$; note that this is also a deterministic commitment. By definition, each MDP $\mathcal{M}_{s}$ has horizon up to $H$ and thus by the induction hypothesis, it admits a mapping $\rho^{\pi|\sigma}(\cdot)$ and a non-negative cost sharing vector $b^{\pi|\sigma}$, satisfying the properties of the lemma. Now, consider any trajectory $\tau\in\Upsilon_{\pi}$ and observe that $\tau=\{\sigma,\tau_{s}\}$ for some $s\in R_{j}$ and some $\tau_{s}\in\Upsilon_{\pi}(s)$. We define
$\rho^{\pi}(\tau):=m_{j}(\max\{g_{j},\rho^{\pi|s}(\tau_{s})\})$
@@
$b^{\pi}_{\sigma\tau}:=\operatorname{E}\left[\rho^{\pi}(\tau)\right]-\operatorname{E}\left[\rho^{\pi|s}(\tau_{s})\right]$
and for all other $s\in S_{\pi}\setminus\{\sigma\}$ and $\tau\in\Upsilon_{\pi}(s)$, we use the same cost share $b^{\pi}_{s\tau}=b^{\pi|s}_{s\tau}$ that was used in $\mathcal{M}_{s}$.
  • Cost sharing: This holds trivially by the definition of the cost shares in the starting state $\sigma$, the fact that we don’t change the cost shares in any other state $s\neq\sigma$, the fact that $v(\tau)=v(\tau_{s})$ for any suffix $\tau_{s}$ of $\tau$ and the induction hypothesis.
  • Cost dominance: By the induction hypothesis and the fact that we maintain the cost shares of every state $s\neq\sigma$, we only need to prove the inequality for $\sigma$. We have
    $\sum_{\tau\in\Upsilon_{\pi}}p_{\pi}(\tau)b^{\pi}_{\sigma\tau}$ $=\sum_{s\in R_{j},\tau_{s}\in\Upsilon_{\pi}(s)}p_{\pi}(\{\sigma,\tau_{s}\})(\operatorname{E}\left[\rho^{\pi}(\{\sigma,\tau_{s}\})\right]-\operatorname{E}\left[\rho^{\pi|s}(\tau_{s})\right])$
    $=\sum_{\tau\in\Upsilon_{\pi}}p^{\pi}(\tau)\operatorname{E}\left[\rho^{\pi}(\tau)\right]-\sum_{s\in R_{j}}\mathcal{D}(\sigma,a_{j},s)\sum_{\tau_{s}\in\Upsilon_{\pi}(s)}p^{\pi|s}(\tau_{s})\operatorname{E}\left[\rho^{\pi|s}(\tau_{s})\right]$
@@
    $\leq\operatorname{E}\left[\hat{Z}_{j}\right]-\operatorname{E}\left[Z_{j}\right]=c(a_{j}).$
    and the statement follows by noting $p^{\pi}(\sigma)=1$. Here the first equation follows from the definition of $b^{\pi}_{\sigma\tau}$ and a decomposition of the trajectories $\tau$; the second just rewrites the terms separately; the third is by the action independence of $\rho^{\pi}$ (proven below), and by the induction hypothesis applied to $\mathcal{M}_{s}$ similarly; the fourth uses property (2) in Lemma 3.6 for the first term and the definition of $Z_{j}$ for the second term; and the last equality follows from the definition of $\hat{Z}_{j}$.
  • Action independence: Drawing a trajectory $\tau\sim p_{\pi}$ is equivalent to first drawing a state $s\in R_{j}$ from the distribution $\mathcal{D}(\sigma,a_{j},\cdot)$ and then drawing a trajectory $\tau_{s}\in\Upsilon_{\pi}(s)$ from $p_{\pi|s}$. Consider drawing $\tau$ in this manner and then sampling from the distribution $\rho^{\pi|s}(\tau_{s})$. By the induction hypothesis and the definition of $Z_{j}$, this provides us with a sample drawn from $Z_{j}$. Then, $\max\{g_{j},\rho^{\pi|s}(\tau_{s})\}$ with $\tau$ drawn in this manner corresponds to an instantiation of $\hat{Z}_{j}$, and $m_{j}$ applied to that instantiation results in an instantiation of $W^{*}_{\mathcal{M}}$ by the definition of $m_{j}$ and property (1) in Lemma 3.6.
@@ 4 Local Approximation and Composition Theorems
Recall that the commitment gap of a CICS instance is defined as the ratio between the cost of the optimal committing policy and the cost of the optimal (non-committing) policy. Let $\mathbb{I}=(\mathbb{M},\mathcal{F})$ be an instance of matroid-min-CICS and let $\pi_{i}$ be a commitment for $\mathcal{M}_{i}\in\mathbb{M}$. We will establish bounds on the commitment gap achieved by $\mathcal{P}=(\pi_{1},\cdots,\pi_{n})$ by quantifying for each $i\in[n]$ the performance of $\pi_{i}$ in the local game $(\mathcal{M}_{i},y)$.
Recall that the performance of an optimal algorithm for the local game $(\mathcal{M},y)$ is described by the optimality curve $f_{\mathcal{M}}(y)$. The performance of the commitment $\pi$ in the same game is given by $f_{\mathcal{M}^{\pi}}(y)$. Scully and Doval (2024) define local approximation by relating these two quantities: [^8] Their definition is provided in the context of optional-inspection PB, and we extend the same definition to general MDPs.
@@
Let $\mathcal{M}$ be any MDP and let $\alpha\geq 1$. We say that a commitment $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ is an $\alpha$-local approximation for $\mathcal{M}$ if $f_{\mathcal{M}^{\pi}}(\alpha y)\leq\alpha\cdot f_{\mathcal{M}}(y)$ for all $y\in\real$.
Restating the inequality in terms of the water-filling costs provides a more intuitive definition as well as a composition theorem. Theorems 3.2 and 3.4 together imply that we can bound the commitment gap as:
$\mathrm{CG}(\mathbb{I})\leq\min_{(\pi_{1},\cdots,\pi_{n})\in\mathcal{C}\left(\mathbb{I}\right)}\frac{\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}^{\pi_{i}}_{i}}\right]}{\operatorname{E}\left[\min_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}\right]}.$
On the other hand, using Definition 7, an $\alpha$-local approximation for $\pi_{i}$ with respect to $\mathcal{M}_{i}$ can be restated as a second-order stochastic dominance condition:
$\operatorname{E}\left[\min\{y,W^{*}_{\mathcal{M}_{i}^{\pi_{i}}}\}\right]\leq\operatorname{E}\left[\min\{y,\alpha W^{*}_{\mathcal{M}_{i}}\}\right]\;\;\forall y\in\real.$ (2)
Combining these inequalities provides the following composition result.
@@
Let $\mathbb{I}=(\mathbb{M},\mathcal{F})$ be any instance of matroid-min-CICS, where each constituent MDP $\mathcal{M}_{i}$ admits an $\alpha$-local approximation under some commitment $\pi_{i}\in\mathcal{C}\left(\mathcal{M}_{i}\right)$. Then, $\mathrm{CG}(\mathbb{I})\leq\alpha$.
The proof of Theorem 4.1 is straightforward from our previous discussion and is presented in Appendix E. The theorem also provides a recipe for designing approximation algorithms for CICS: if we can efficiently identify a commitment $\pi_{i}$ for each MDP $\mathcal{M}_{i}$ that achieves an $\alpha$-local approximation, then we can simply restrict each MDP $\mathcal{M}_{i}$ to the corresponding Markov chain $\mathcal{M}_{i}^{\pi_{i}}$ and run the (efficient) water filling index policy; the resulting committing policy is guaranteed to be an $\alpha$-approximation to the unrestricted optimum.
Geometric Intuition for Local Approximation.
It is worth emphasizing that $\alpha$-local approximation is not equivalent to simply achieving a standard $\alpha$-approximation in the local game. Slightly rearranging the condition in Definition 8, we have that for commitment $\pi$ to be an $\alpha$-local approximation, we require for all $y\in\mathbb{R}$,
$f_{\mathcal{M}^{\pi}}(y)\leq\alpha f_{\mathcal{M}}(y/\alpha).$
In contrast, for commitment $\pi$ to be a standard $\alpha$-approximation for the local game, we require only
$f_{\mathcal{M}^{\pi}}(y)\leq\alpha f_{\mathcal{M}}(y),$
the difference being $y$ instead of $y/\alpha$ on the right-hand side. As illustrated in Figure 4, we can understand both of the above conditions as saying that the graph of $f_{\mathcal{M}^{\pi}}(y)$ needs to lie below a geometrically scaled version of the graph of $f_{\mathcal{M}}(y)$. However, the scaling used is different for the two conditions.
  1. For standard $\alpha$-approximation, we vertically scale up $f_{\mathcal{M}}(y)$ by a factor of $\alpha$.
  2. For $\alpha$-local approximation, we vertically and horizontally scale up $f_{\mathcal{M}}(y)$ by a factor of $\alpha$.
The additional horizontal scaling in local approximation is critical in proving 2 and thereby obtaining Theorem 4.1. Standard $\alpha$-approximation implies only a weaker version of 2 with the $y$ on the right-hand side replaced by $\alpha y$. Working through the proof of Theorem 4.1 with this change, one would eventually obtain a commitment gap bound of $\alpha^{n}$, scaling exponentially [^9] Appendix C exhibits an explicit example of this exponential scaling. in the number of MDPs $n$. Using the right local approximation notion is thus critical for obtaining approximation guarantees that do not depend on $n$.
[…]
Figure 4: For an MDP $\mathcal{M}$ with a given optimality curve (left), we contrast two approximation notions: standard $\alpha$-approximation (center) vs. $\alpha$-local approximation (right). In both cases, achieving the approximation requires designing a commitment whose optimality curve lies below a scaled version of the optimal cost (purple curves). For standard $\alpha$-approximation, the curve is scaled up vertically by a factor of $\alpha$. For $\alpha$-local approximation, the curve is scaled up diagonally, i.e. both vertically and horizontally, by a factor of $\alpha$.
Establishing Local Approximation.
We note that in some cases, the surrogate costs of an MDP might not admit a simple structure, making second-order stochastic dominance conditions difficult to establish. We can alternatively establish local approximation through a stronger but conceptually easier relationship between the surrogate costs: namely, that $\alpha W^{*}_{\mathcal{M}}$ first-order stochastically dominates $W^{*}_{\mathcal{M}_{\pi}}$. Formally, we define the following alternate notion of approximation, where for a quantile $q\in[0,1]$ and a random variable $X$, we let $X(q)$ denote the $q$th quantile value of the r.v.: $X(q)=\inf\{x:Pr[X\leq x]\geq q\}$.
Definition 9 (Pointwise Approximation).
@@
Let $\mathcal{M}$ be any MDP and let $\alpha\geq 1$. We say that a commitment $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ is an $\alpha$-pointwise approximation for $\mathcal{M}$ if $W^{*}_{\mathcal{M}^{\pi}}(q)\leq\alpha\cdot W^{*}_{\mathcal{M}}(q)$ for all $q\in[0,1]$.
The following is immediate from the fact that first-order stochastic dominance implies second-order stochastic dominance, but the converse need not hold.
Fact 4.2.
If $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ is an $\alpha$-pointwise approximation for $\mathcal{M}$, it also an $\alpha$-local approximation for $\mathcal{M}$.
Finally, we note that by definition we have $f_{\mathcal{M}}(y)=\min_{\pi\in\mathcal{C}\left(\mathcal{M}\right)}f_{\mathcal{M}^{\pi}}(y)$. Moreover, because randomized commitments generate optimality curves as averages over those of deterministic commitments, we can write $f_{\mathcal{M}}(y)$ as the minimum over the optimality curves of deterministic commitments. Thus, we can obtain local or pointwise approximation guarantees by showing that the commitment $\pi$ satisfies the desired definition against all other deterministic commitments $\pi^{\prime}\in\mathcal{C}\left(\mathcal{M}\right)$. This simple observation can prove essential when establishing these conditions, as the water-filling surrogate costs of Markov chains have a much more intuitive structure (Lemma 3.1) than the surrogate costs of MDPs (Lemma 3.5). For example, while the surrogate costs of the MDPs corresponding to Additive PBs (Section 6) are very complicated, the surrogate costs of committing policies for the same setting admit an exceptionally simple form that can be exploited to prove a pointwise approximation.
Beyond Local Approximation.
For one of our applications, namely the maximization version of Pandora’s box with optional inspection (Section 8), local approximation turns out to be too stringent of a condition. As such, we introduce a slight weakening of local approximation, which we call semilocal approximation (Definition 12), that is tailored to this application. Roughly speaking, local approximation requires purely multiplicative suboptimality, while semilocal approximation allows for some additive suboptimality, too. See Section 8 for details.